# Ensemble-Based Data Assimilation for Material Model Characterization in High-Velocity Impact

### v1

@@ Abstract
High-fidelity simulations are essential for understanding and predicting the behavior of materials under high-velocity impact (HVI) in both fundamental research and practical applications. However, their accuracy relies on material models and parameters that are traditionally obtained through manual fitting to multiple time- and labor-intensive experiments. In this study, we develop an ensemble-based data assimilation (DA) framework that automatically and simultaneously calibrates plasticity, fracture, and equation of state (EOS) model parameters in HVI simulations using measured data from a single HVI test. The framework integrates Smoothed Particle Hydrodynamics for HVI simulations, the ensemble Kalman filter (EnKF) for parameter refinement, and adaptive covariance inflation to mitigate underestimation of uncertainty. We demonstrate the approach using synthetic back-face deflection data of an AZ31B magnesium plate to identify parameters in the Johnson-Cook plasticity and fracture models and the Mie-Grüneisen EOS. Test cases include under- and over-biased initial guesses, as well as limited observational data. The results show that, when the observational data are sufficiently sensitive to the parameters, the EnKF-based framework can accurately and efficiently recover the material parameters in five iterations with convergent ensemble standard deviation. In contrast, insensitive parameters tend to converge to incorrect values and are characterized by persistently large ensemble standard deviations. Furthermore, studies on observation quantity reveal that the limited data can still achieve convergence to the true values but with more iterations. Under extreme prior bias, sensitive parameters can exhibit a drift-then-stall behavior in which posterior spreads shrink and prediction errors decrease, yet small residual biases persist. In practical applications where the true parameters are unknown, the ensemble standard deviation thus provides a diagnostic tool to assess parameter sensitivity and calibration accuracy. Overall, this study demonstrates the robustness and efficiency of the proposed DA framework for material model characterization in HVI problems.
@@ 1 Introduction
The remainder of the paper is organized as follows. Section 2 formulates the HVI model-calibration problem. Section 3 outlines the fundamentals of the SPH numerical method and presents the governing equations for material modeling. Section 4 introduces the EnKF equations and describes the development of the EnKF framework. Section 5 details the numerical experiments, including the specification of model inputs and outputs, sensitivity screening, generation of synthetic observations, and presentation of calibration results. Section 6 discusses the main findings, and Section 7 concludes the paper with a concise summary.
@@ 4.1 Ensemble Kalman Filter
$\bm{x}_{l+1}^{(i)}=\hat{\bm{x}}_{l+1}^{(i)}+\mathbf{K}_{l}\left[\bm{y}_{\rm{obs}}^{(i)}-\mathbf{H}\hat{\bm{x}}_{l+1}^{(i)}\right],\quad i=1,\dots,N_{\rm{e}},$ (21)
@@
$\mathbf{K}_{l}=\mathbf{P}_{l}\mathbf{H}^{\text{T}}\left[\mathbf{H}\mathbf{P}_{l}\mathbf{H}^{\text{T}}+\mathbf{R}\right]^{-1},$ (22)
and $\bm{y}_{\rm{obs}}^{(i)}$ is the perturbed observation.
@@ 4.2 Covariance Inflation
$\beta_{l}^{(j)}=1+\kappa\,\frac{\sqrt{\mathbf{P}_{l-1}(j,j)}-\sqrt{\mathbf{P}_{l}(j,j)}}{\sqrt{\mathbf{P}_{l}(j,j)}},\quad j=1,\dots,N_{y}+N_{u},$ (24)
where $\beta_{l}^{(j)}$ is the $j$-th component of the multiplicative inflation vector $\bm{\beta}_{l}$, and $\kappa$ is a scalar. $\mathbf{P}_{l}(j,j)$ denotes the $j$-th diagonal entry of the covariance matrix $\mathbf{P}_{l}$. Thus, $\sqrt{\mathbf{P}_{l-1}(j,j)}$ and $\sqrt{\mathbf{P}_{l}(j,j)}$ correspond to the standard deviations of the prior and posterior ensembles, respectively, for the $j$-th state component at iteration $l$.
@@ 4.3 EnKF Inversion Framework
2Initialization:;
3
4for $i=1,\dots,N_{\text{e}}$ do
@@ 11 Forecast:;
16   end
17
18  Form the augmented state $\bm{x}_{l}$ (i.e., $\hat{\bm{x}}_{l+1}$) using Eq. (16);
19
20 Analysis:;
21
@@ 20 Analysis:;
27
28 Covariance inflation:;
29
30  Compute ensemble covariance $\mathbf{P}_{l+1}$ using Eq. (19);
31
32  Compute the multiplicative factor $\beta_{l+1}^{(j)},j=1,\dots,N_{y}+N_{u}$ using Eq. (24);
@@ 28 Covariance inflation:;
5 Numerical Experiments
@@ 5.2 HVI Behavior
In this subsection, we evaluate the HVI response of the plate using the baseline material parameters listed in Tables 1-3. The objective is to assess the fidelity and predictive capability of the SPH method. Fig. 2 presents the level contours of the maximum principal stress and temperature at three representative time instants: $t=3.0~\upmu\text{s}$, $12.0~\upmu\text{s}$, and $20.0~\upmu\text{s}$. For clarity of presentation, the projectile and half of the target plate are removed during post-processing to enable an unobstructed view of field evolution within the contact region.
@@ 5.3 Synthetic Observations
The synthetic perturbed observation $\bm{y}_{\rm obs}^{(i)}$ is generated according to
$\bm{y}_{\rm obs}^{(i)}=\bm{y}_{\rm true}+\bm{\eta}_{0}+\bm{\eta}^{(i)},$ (25)
where $\bm{y}_{\rm true}$ represents the model output at the baseline parameter values. The term $\bm{\eta}_{0}$ denotes a systematic bias, set to 0.001-0.001 mm, while $\bm{\eta}^{(i)}$ is a zero-mean Gaussian random error, i.e. $\bm{\eta}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$. The superscript $(i)$ implies that the observation error is sampled in each ensemble. The observation covariance $\mathbf{R}$ is assumed to be diagonal, and all measured deflections share the same standard deviation. This standard deviation is assumed to be a fixed value, which is estimated by multiplying the reported random error $4\%$ of 3D-DIC by the smallest observable displacement listed in Table 4, yielding a value of $0.01$ mm. This procedure produces a dataset that captures both the spatio-temporal distribution and noise characteristics of typical 3D-DIC measurements, thereby providing a realistic testing environment for evaluating the EnKF framework.
@@ 5.4 Input Space Reduction
Figure 5: Back-face deflection profiles produced by the extreme perturbations $\theta_{i}^{-}=0.75\,\theta_{i}$ (dashed) and $\theta_{i}^{+}=1.25\,\theta_{i}$ (solid) of the retained parameters: a) $C$, (b) $D_{4}$, and (c) $\gamma_{0}$. Curves are colored by time, red: $t=10.0\,\upmu\rm{s}$; green: $t=11.0\,\upmu\rm{s}$; blue: $t=12.0\,\upmu\rm{s}$.
@@ 5.5 EnKF Inversion Results
We employ the developed EnKF inversion framework to calibrate the model parameters $C$, $D_{4}$, and $\gamma_{0}$. The ensemble size is set to $N_{\rm e}=100$ and the total number of EnKF iterations to $M=20$. This ensemble size provides sufficient accuracy while maintaining a reasonable computational cost, and the chosen number of iterations ensures the convergence of the analysis. The initial guess for the unknown parameters is taken as the mean of the initial ensemble $\bm{u}_{0}$. The prior covariance matrix $\mathbf{C}_{0}$ is assumed diagonal, with the standard deviation of each component set to $10\%$ of its initial guess.
@@
    Case 4: strongly biased initial guess, $\bm{u}_{0}=2.5\,\bm{u}_{\text{true}}=[0.0325,\,1.1845,\,3.85]^{\text{T}}$, where the initial guesses for all parameters deviate by $150\%$ from their true values;
The EnKF results are plotted in Figs. 6-8. All figures share a unified layout: subfigures (a)-(c) display the iteration trajectories of $C$, $D_{4}$, and $\gamma_{0}$, respectively, where the red dashed line marks the truth, the blue dots represent the ensemble mean, and the light-gray shading denotes the min-max ensemble range. Subfigure (d) reports the time-resolved RMSE of the simulated back-face deflection before and after assimilation, both computed against the reference displacement at the baseline values of the unknown parameters. In addition, a concise comparison of the ensemble mean, standard deviation (std), and relative error for all four cases is summarized in Table 5.
@@
Figure 6: EnKF results for Case 1 with initial ensemble mean $\bm{u}_{0}=0.75\,\bm{u}_{\text{true}}$ ($25\%$ bias). (a-c): Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d): Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
@@
Figure 7: EnKF results for Case 2 with initial ensemble mean $\bm{u}_{0}=1.25\,\bm{u}_{\text{true}}$ ($25\%$ bias). (a-c): Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d): Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
@@
Figure 8: EnKF results for Case 3 with initial ensemble mean $\bm{u}_{0}=0.75\,\bm{u}_{\text{true}}$, with $N_{\rm{o}}=10$ at each observation time. (a-c): Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d): Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
To stress-test the robustness of the EnKF inversion framework, we significantly bias the initial guess in Case $4$ such that the true parameters lie outside the initial ensemble spread. Fig. 9 summarizes the results. Even under this extreme misspecification, the filter remains numerically stable and effectively reduces predictive error and posterior uncertainty within the impact window. The two sensitivity-dominant parameters, $C$ and $\gamma_{0}$, show a clear trend toward the truth during the early iterations and then plateau with a persistent positive bias. Consistent with Table 5, initial discrepancies of $149.50\%$ for $C$ and $149.38\%$ for $\gamma_{0}$ decrease to $57.63\%$ and $16.50\%$, respectively. In contrast, $D_{4}$ is poorly identifiable with the present data and initialization and exhibits a wide credible band, ending with a residual bias of $126.43\%$. As illustrated in Fig. 9(d), the post-calibration RMSE consistently demonstrates a reduction compared to the baseline, the most notable improvements occurring after $t\approx 3~\upmu\rm{s}$, thus indicating the improvement of the EnKF inversion process.
Figure 9: EnKF results for Case 4 with initial ensemble mean $\bm{u}_{0}=[0.0325,1.1845,3.85]^{\text{T}}$. (a-c): Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d): Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.

### v2

@@ Abstract
High-fidelity simulations are essential for understanding and predicting the behavior of materials under high-velocity impact (HVI) in both fundamental research and practical applications. However, their accuracy relies on material models and parameters that are traditionally obtained through manual fitting to multiple time- and labor-intensive experiments. In this study, we develop an ensemble-based data assimilation (DA) framework that automatically and simultaneously calibrates selected plasticity, fracture, and equation of state (EOS) model parameters in HVI simulations using measured data from a single HVI test. The framework integrates Smoothed Particle Hydrodynamics for HVI simulations, the ensemble Kalman filter (EnKF) for parameter refinement, and adaptive covariance inflation to mitigate underestimation of uncertainty. We first show that, for a simple test problem, the framework achieves a computational efficiency at least one order of magnitude higher than traditional Markov chain Monte Carlo methods at comparable identification accuracy. We then demonstrate the approach using synthetic back-face deflection data of an AZ31B magnesium plate to identify representative parameters in the Johnson-Cook plasticity and fracture models and the Mie-Grüneisen EOS. Test cases include under- and over-biased initial guesses, as well as limited observational data. The results show that, when the observational data are sufficiently sensitive to the parameters, the EnKF-based framework can accurately and efficiently recover the material parameters in five iterations with convergent ensemble standard deviation. In contrast, insensitive parameters tend to converge to incorrect values and are characterized by persistently large ensemble standard deviations. Furthermore, studies on observation quantity reveal that the limited data can still achieve convergence to the true values but with more iterations. Under extreme prior bias, the proposed parameter rejuvenation strategy effectively ensures that sensitive parameters migrate toward the true values even when the truth lies outside the initial ensemble spread. In practical applications where the true parameters are unknown, the ensemble standard deviation thus provides a diagnostic tool to assess parameter sensitivity, identifiability, and potential non-uniqueness. Overall, this study demonstrates the robustness and efficiency of the proposed DA framework for material model characterization in HVI problems.
@@ 1 Introduction
The remainder of the paper is organized as follows. Section 2 formulates the HVI model-calibration problem. Section 3 outlines the fundamentals of the SPH numerical method and presents the governing equations for material modeling. Section 4 introduces the EnKF equations and describes the development of the EnKF framework. Section 5 compares this framework with a traditional MCMC method using a ring-down problem. Section 6 details the numerical experiments, including the specification of model inputs and outputs, sensitivity screening, generation of synthetic observations, and presentation of calibration results. Section 7 discusses the main findings, and Section 8 concludes the paper with a concise summary.
@@ 4.1 Ensemble Kalman Filter
$\bm{x}_{l+1}^{(i)}=\hat{\bm{x}}_{l+1}^{(i)}+\mathbf{K}_{l}\left[\bm{y}_{\rm{pert}}^{(i)}-\mathbf{H}\hat{\bm{x}}_{l+1}^{(i)}\right],\quad i=1,\dots,N_{\rm{e}},$ (21)
@@
$\mathbf{K}_{l}=\mathbf{P}_{l}\mathbf{H}^{\text{T}}\left[\mathbf{H}\mathbf{P}_{l}\mathbf{H}^{\text{T}}+\mathbf{R}\right]^{-1}.$ (22)
In Eq. (21), $\bm{y}_{\rm pert}^{(i)}$ denotes the perturbed observation
$\bm{y}_{\rm pert}^{(i)}=\bm{y}_{\rm obs}+\bm{\eta}^{(i)},$ (23)
where $\bm{y}_{\rm obs}$ is the observation vector and $\bm{\eta}^{(i)}$ is a zero-mean Gaussian perturbation, i.e., $\bm{\eta}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$. The superscript $(i)$ implies that the observation error is sampled in each ensemble.
@@ 4.2 Covariance Inflation
$\bm{\beta}_{l}{(j)}=1+\kappa\,\frac{\sqrt{\mathbf{P}_{l-1}(j,j)}-\sqrt{\mathbf{P}_{l}(j,j)}}{\sqrt{\mathbf{P}_{l}(j,j)}},\quad j=1,\dots,N_{y}+N_{u},$ (25)
where $\bm{\beta}_{l}{(j)}$ is the $j$-th component of the multiplicative inflation vector $\bm{\beta}_{l}$, and $\kappa$ is a scalar. $\mathbf{P}_{l}(j,j)$ denotes the $j$-th diagonal entry of the covariance matrix $\mathbf{P}_{l}$. Thus, $\sqrt{\mathbf{P}_{l-1}(j,j)}$ and $\sqrt{\mathbf{P}_{l}(j,j)}$ correspond to the standard deviations of the prior and posterior ensembles, respectively, for the $j$-th state component at iteration $l$.
@@
While the RTPS method effectively maintains ensemble spread in nominal conditions, extreme initial biases can still lead to filter divergence, where the ensemble variance collapses before the mean migrates to the true solution. To further mitigate this failure mode, we introduce a strategy that selectively re-inflates the parameter subspace toward its initial relative ensemble spread when the filter exhibits concurrent signs of inconsistency and ensemble collapse. We refer to this strategy as “parameter rejuvenation”. Specifically, at iteration $l$ we monitor the normalized misfit
$\mathcal{J}_{l}=\frac{1}{N_{y}}\big(\bm{y}_{\text{obs}}-\mathbf{H}\bar{\bm{x}}_{l}\big)^{\text{T}}\mathbf{R}^{-1}\big(\bm{y}_{\text{obs}}-\mathbf{H}\bar{\bm{x}}_{l}\big),$ (26)
and the coefficient-of-variation (CV) vector of the unknown parameters, whose $j$-th component is defined as
$\bm{c}_{l}{(j)}=\frac{\sqrt{\mathbf{P}_{l}(j+N_{y},j+N_{y})}}{|\overline{\bm{x}}_{l}(j+N_{y})|},$ (27)
where $\sqrt{\mathbf{P}_{l}(j+N_{y},j+N_{y})}$ and $\overline{\bm{x}}_{l}(j+N_{y})$ correspond to the standard deviation and mean of the $j$-th parameter. Rejuvenation is activated only when all three conditions below are satisfied simultaneously:
(1) significant misfit: the predicted output deviates substantially from the observations
$\mathcal{J}_{l}>\mathcal{J}_{\text{t}},\quad\mathcal{J}_{\text{t}}=1.5;$ (28)
(2) stagnation: the relative decrease of misfit over a short window is small
$\frac{\mathcal{J}_{l-w}-\mathcal{J}_{l}}{\mathcal{J}_{l-w}}<\epsilon_{\text{d}},\quad w=3,\ \epsilon_{\text{d}}=0.1;$ (29)
(3) ensemble spread collapse: at least one calibrated parameter has collapsed CV
$\exists j:\quad\bm{c}_{l}{(j)}<\eta_{\text{c}}\,\bm{c}_{0}{(j)},\quad\eta_{\text{c}}=0.5.$ (30)
When parameter rejuvenation is triggered, we bypass RTPS and instead apply a multiplicative re-inflation to the parameter ensemble
$\bm{u}_{l}^{(i)}\leftarrow\bar{\bm{u}}_{l}+\alpha_{\mathrm{r}}\,\big(\bm{c}_{0}\oslash\bm{c}_{l}\big)\odot\big(\bm{u}_{l}^{(i)}-\bar{\bm{u}}_{l}\big),\qquad\alpha_{\mathrm{r}}=0.95,\qquad i=1,\ldots,N_{\mathrm{e}},$ (31)
[…]
@@ 4.3 EnKF Inversion Framework
2Initialization:
3for $i=1,\dots,N_{\text{e}}$ do
@@ 9 Forecast:
16 Analysis:
@@
20
21  Update the state $\bm{x}_{l+1}^{(i)},i=1,\dots,N_{\text{e}}$ using Eq. (21);
22
23 Covariance inflation:
24  Compute ensemble statistics $\overline{\bm{x}}_{l+1}$ and $\mathbf{P}_{l+1}$ using Eqs. (18) and (19);
@@ 23 Covariance inflation:
26  Compute normalized misfit $\mathcal{J}_{l}$ using Eq. (26);
27  Compute coefficient-of-variation $\bm{c}_{l}$ using Eq. (27);
28 if Eqs. (28)-(30) are satisfied then
29      Update the parameters $\bm{u}_{l+1}^{(i)},i=1,\dots,N_{\text{e}}$ using Eq. (31);
30 else
31      Compute multiplicative factor $\beta_{l+1}^{(j)},j=1,\dots,N_{y}+N_{u}$ using Eq. (25);
@@
5 Comparison with a Traditional Bayesian Method
We first compare the EnKF-based approach with a traditional sampling-based MCMC method. In general, MCMC requires on the order of $10^{4}$ (or more) sequential samples to obtain a well-resolved posterior, making it computationally prohibitive for HVI problems in which each forward SPH simulation is expensive. To provide a controlled comparison, we consider an impact-inspired ring-down problem with an analytical forward mapping. The ring-down problem describes the post-impact free vibration of a structural component (e.g., a plate) after the short-duration contact phase of a high-velocity impact has ended. We assume the ensuing response is dominated by the first bending mode and model the back-face deflection using a separable spatio-temporal representation. Because our goal is to calibrate model parameters using back-face deflection data, we define the forward operator $\mathcal{G}(\bm{u})$ mapping the unknown parameters to the back-face deflection as
$\mathcal{G}(\bm{u})=\phi(x)\,q(\bm{u};t),$ (32)
where $x\in[0,L]$ denotes a coordinate along a representative centerline. The spatial variation is represented by a simple symmetric mode shape
$\phi(x)=\sin\!\left(\frac{\pi x}{L}\right).$ (33)
The temporal response follows an underdamped single-degree-of-freedom (SDOF) model
$q(\bm{u};t)=a\exp(-\zeta\omega_{\text{n}}t)\,\sin(\omega_{\text{n}}\sqrt{1-\zeta^{2}}t),$ (34)
parameterized by the unknown vector to be calibrated
$\bm{u}=[a,\ \omega_{\text{n}},\ \zeta]^{\text{T}},$ (35)
where $a$ is the amplitude associated with impact severity, $\omega_{\text{n}}$ is the natural frequency governed by the effective stiffness-to-mass ratio, and $\zeta$ is the damping ratio capturing energy dissipation. This model yields an exponentially decaying sinusoidal deflection field that captures the essential ring-down behavior while avoiding explicit modeling of the contact phase and nonlinear impact dynamics.
The observation data consist of displacement measurements collected at $20$ equally spaced locations along the centerline (with $L=1$) and at three time instants $\{6\times 10^{-4},\,1.0\times 10^{-3},\,1.4\times 10^{-3}\}\,$s, yielding a stacked observation dimension of $N_{y}=60$. The observations are generated using the true parameters $\bm{u}_{\rm true}=\big[3.3~\text{mm},\ 2200~\text{rad}/\text{s},\ 0.06\big]^{\text{T}},$ and corrupted with additive Gaussian noise of zero mean and standard deviation $\sigma_{\eta}=1.0\times 10^{-2}$ mm, i.e., $\bm{y}_{\rm obs}=\bm{y}_{\rm ref}+\bm{\eta}$ with $\bm{\eta}\sim\mathcal{N}(\bm{0},\sigma_{\eta}^{2}\bm{I})$. For the EnKF simulations, we use an ensemble size of $N_{e}=100$ and run $M=100$ iterations. We also solve this example using a Random-Walk Metropolis MCMC (RWM-MCMC) scheme by running $3\times 10^{4}$ steps and discarding the first $5\times 10^{3}$ samples as burn-in; posterior samples are recorded every $10$ steps (thinning factor $10$). Both EnKF and RWM-MCMC are initialized using the same Gaussian prior in transformed coordinates $\bm{\theta}=[\log(a),\,\log(\omega_{n}),\,\mathrm{logit}(\zeta)]^{\mathrm{T}}$, where $\mathrm{logit}(\zeta)=\log\!\big(\zeta/(1-\zeta)\big)$. Specifically, $\bm{\theta}\sim\mathcal{N}(\bm{\theta}_{0},\mathrm{diag}(\bm{\sigma}_{\theta}^{2}))$ with $\bm{\theta}_{0}=[\log(2.0),\,\log(1800),\,\log(0.05/0.95)]^{\mathrm{T}}$ and $\bm{\sigma}_{\theta}=[0.35,\,0.25,\,0.60]^{\mathrm{T}}$.
Fig. 1(a) compares the evolution of the root-mean-square error (RMSE) of the back-face displacement at the selected locations and time instants as a function of the number of forward evaluations. A more detailed comparison of computational cost is provided in Table 1. Because burn-in contributes to the wall-clock cost of sequential MCMC, it is included when comparing the computational cost of RWM-MCMC with EnKF. As shown, the EnKF converges within the first $4$ iterations, requiring $500$ forward evaluations, whereas RWM-MCMC requires approximately $6000$ iterations (and forward evaluations) to reach a comparable level of accuracy. The final RMSE obtained by EnKF is $1.86\times 10^{-3}$ mm, which is comparable to that of RWM-MCMC ($1.95\times 10^{-3}$ mm), indicating that both methods recover posterior means that are close to the true values. To quantify computational efficiency in a fair and interpretable manner, we introduce a Cost-to-Misfit metric, $\mathcal{C}$, defined as the number of forward evaluations required to reduce the RMSE below a prescribed threshold. For example, for a target RMSE of $2.5\times 10^{-3}$ mm (Fig. 1(a)), RWM-MCMC requires $5512$ forward evaluations, while EnKF requires only $400$, i.e., an approximately $13.8\times$ reduction in forward evaluations (at least one order of magnitude).
Figure 1: Comparison of EnKF and RWM-MCMC for the ring-down model: (a) RMSE of the back-face displacement versus the number of forward evaluations. The dashed line indicates an illustrative target threshold $\mathrm{RMSE}$. (b)-(d) Pairwise posterior projections of the inferred parameter distributions. The EnKF results are shown as Gaussian-fit probability-mass contours at levels $\{0.10,0.20,0.30,0.40,0.50,0.60,0.70,0.80,0.90,0.99\}$.
Table 1: Efficiency comparison on the ring-down model.
Method Type Total Evals. $\mathcal{C}(\mathrm{RMSE}_{\rm target})$ Final RMSE (mm)
EnKF Parallel 10,100 400 $1.86\times 10^{-3}$
RWM-MCMC Sequential 30,000 5,512 $1.95\times 10^{-3}$
Figs. 1(b)-(d) compare the pairwise posterior distributions obtained from EnKF with those from RWM-MCMC. The two sets of posteriors are in close agreement, indicating that EnKF recovers a posterior distribution comparable to that of RWM-MCMC, but at a substantially lower computational cost.
6 Application to HVI
@@ 6.2 HVI Behavior
In this subsection, we evaluate the HVI response of the plate using the baseline material parameters listed in Tables 2-4. The objective is to assess the fidelity and predictive capability of the SPH method. Fig. 3 presents the level contours of the maximum principal stress and temperature at three representative time instants: $t=3.0~\upmu\text{s}$, $12.0~\upmu\text{s}$, and $20.0~\upmu\text{s}$. For clarity of presentation, the projectile and half of the target plate are removed during post-processing to enable an unobstructed view of field evolution within the contact region.
@@ 6.3 Synthetic Observations
The synthetic observation $\bm{y}_{\rm obs}$ is generated according to
$\bm{y}_{\rm obs}=\bm{y}_{\rm true}+\bm{\eta}_{0},$ (36)
where $\bm{y}_{\rm true}$ represents the model output at the baseline parameter values. The term $\bm{\eta}_{0}$ denotes a systematic bias, set to 0.001-0.001 mm. The observation covariance $\mathbf{R}$ is assumed to be diagonal, and all measured deflections share the same standard deviation. This standard deviation is assumed to be a fixed value, which is estimated by multiplying the reported random error $4\%$ of 3D-DIC by the smallest observable displacement listed in Table 5, yielding a value of $0.01$ mm. This procedure produces a dataset that captures both the spatio-temporal distribution and noise characteristics of typical 3D-DIC measurements, thereby providing a realistic testing environment for evaluating the EnKF framework.
@@ 6.4 Input Space Reduction
Figure 6: Back-face deflection profiles produced by the extreme perturbations $\theta_{i}^{-}=0.75\,\theta_{i}$ (dashed) and $\theta_{i}^{+}=1.25\,\theta_{i}$ (solid) of the retained parameters: (a) $C$, (b) $D_{4}$, and (c) $\gamma_{0}$. Curves are colored by time, red: $t=10.0\,\upmu\rm{s}$; green: $t=11.0\,\upmu\rm{s}$; blue: $t=12.0\,\upmu\rm{s}$.
@@ 6.5 EnKF Inversion Results
We employ the developed EnKF inversion framework to calibrate the model parameters $C$, $D_{4}$, and $\gamma_{0}$. The initial guess for the unknown parameters is taken as the mean of the initial ensemble $\bm{u}_{0}$. The prior covariance matrix $\mathbf{C}_{0}$ is assumed diagonal, with the standard deviation of each component set to $10\%$ of its initial guess.
@@
    Case 4: strongly biased initial guess, $\bm{u}_{0}=2.5\,\bm{u}_{\text{true}}$, where the initial guesses for all parameters deviate by $150\%$ from their true values;
Before presenting the detailed calibration results, we first justify the choice of ensemble size $N_{\text{e}}$ and the number of EnKF iterations $M$ using the representative setup of Case 1. To assess convergence, we compute at each EnKF iteration the RMSE of the back-face deflection relative to the ground-truth response and the relative error of the calibrated parameters with respect to their true values. Fig. 7 summarizes the results for three ensemble sizes, $N_{\text{e}}\in\{20,50,100\}$, with all other settings identical to Case 1. As shown, smaller ensembles lead to noticeably larger errors in both the back-face deflection and the inferred parameters. In particular, the RMSE of the back-face deflection is approximately $4.17\times 10^{-3}$ mm for $N_{\text{e}}=20$, whereas it decreases to $7.77\times 10^{-4}$ mm for $N_{\text{e}}=100$. This improvement is expected, as larger ensembles provide a more accurate representation of the forecast covariance and more adequate exploration of the parameter space. Accordingly, we use $N_{\mathrm{e}}=100$ in the following studies. With this choice, the EnKF converges rapidly, reaching a stable solution within roughly the first 55-88 iterations. Although most of the data-fit improvement occurs early, we adopt $M=20$ as a conservative upper bound to ensure robustness in cases with large initial biases (e.g., Case 4).
Figure 7: Convergence histories for ensemble sizes $N_{\mathrm{e}}\in\{20,50,100\}$ using the Case 1 setup: (a) RMSE of the back-face deflection. (b)-(d) Relative errors of the posterior-mean estimates for the calibrated parameters $C$, $D_{4}$, and $\gamma_{0}$, respectively.
The EnKF results for the four cases are plotted in Figs. 8-11. All figures share a unified layout: subfigures (a)-(c) display the iteration trajectories of $C$, $D_{4}$, and $\gamma_{0}$, respectively, where the red dashed line marks the truth, the blue dots represent the ensemble mean, and the light-gray shading denotes the min-max ensemble range. Subfigure (d) reports the time-resolved RMSE of the simulated back-face deflection before and after assimilation, both computed against the reference displacement at the baseline values of the unknown parameters. In addition, a concise comparison of the ensemble mean, standard deviation (std), and relative error for all four cases is summarized in Table 6.
@@
Figure 8: EnKF results for Case 1 with initial ensemble mean $\bm{u}_{0}=0.75\,\bm{u}_{\text{true}}$ ($25\%$ bias): (a)-(c) Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d) Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
@@
Figure 9: EnKF results for Case 2 with initial ensemble mean $\bm{u}_{0}=1.25\,\bm{u}_{\text{true}}$ ($25\%$ bias): (a)-(c) Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d) Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
@@
Figure 10: EnKF results for Case 3 with initial ensemble mean $\bm{u}_{0}=0.75\,\bm{u}_{\text{true}}$, with $N_{\rm{o}}=10$ at each observation time: (a)-(c) Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d) Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.
To stress-test the robustness of the EnKF inversion framework, we significantly bias the initial guess in Case $4$ such that all the true parameters lie outside the initial ensemble spread. Fig. 11 summarizes the results. Even under this extreme misspecification, the filter remains numerically stable and effectively reduces predictive error and posterior uncertainty within the impact window. The two sensitive parameters, $C$ and $\gamma_{0}$, converge to their true values after approximately $15$ iterations. As reported in Table 6, the initial discrepancies of about $150\%$ for both parameters are reduced to $0.81\%$ and $2.16\%$, respectively, while their ensemble spreads decrease by roughly one order of magnitude. In contrast, the insensitive parameter $D_{4}$ remains biased ($136.94\%$) and retains a relatively large ensemble spread, indicating that the posterior uncertainty appropriately reflects limited identifiability rather than exhibiting false confidence. Moreover, Fig. 11(d) shows that the post-calibration RMSE decreases markedly, confirming that the calibrated model reproduces the observed response even under a severely biased initial guess.
It is worth noting that the proposed parameter rejuvenation strategy is not activated in Cases 11-33, because the sensitive parameters $C$ and $\gamma_{0}$ converge very quickly to their true values and the insensitive parameter $D_{4}$ retains a relatively large ensemble spread. However, this strategy plays a critical role in Case $4$, where a strongly biased initial guess is used. In particular, Figs. 11(a) and (c) show that parameter rejuvenation inflates the parameter covariance between iterations $5$ and $15$. Following this inflation, the parameter estimates migrate toward the true values. We further examine a baseline case with the same setup as Case $4$ but without parameter rejuvenation, as detailed in Appendix A. In this baseline case, the ensemble variance collapses rapidly, causing the filter to lock into an incorrect solution before the mean can migrate toward the true parameter values.
Figure 11: EnKF results for Case 4 with initial ensemble mean $\bm{u}_{0}=2.5\,\bm{u}_{\text{true}}$ ($150\%$ bias): (a)-(c) Iteration histories of inferred parameters $C$, $D_{4}$, and $\gamma_{0}$. (d) Time-resolved RMSE of the back-face deflection before and after assimilation with respect to the true deflection over the 0-$12\,\upmu\rm{s}$ observation window.